# Dynamic Open-Vocabulary 3D Scene Graphs for Long-term Language-Guided Mobile Manipulation

### v4

@@ I Introduction
Mobile robot manipulation in real-world environments is increasingly expected to perform complex, long-term tasks that require adaptability and resilience to changes. These dynamic environments are characterized by frequent alterations caused by human activities, robot interactions, or the inherent variability of the surroundings. Traditional robotic systems often fall short in such settings because they typically rely on the assumption of a static or minimally changing environment [1, 2, 3, 4, 5]. This limitation restricts their applicability in real-world scenarios where adaptability is crucial.
In this work, we enhance robotic capabilities by introducing a novel and practical robotic framework, the DovSG system. This framework comprises five key modules: perception, memory, task planning, navigation, and manipulation, as illustrated in Fig. Dynamic Open-Vocabulary 3D Scene Graphs for Long-term Language-Guided Mobile Manipulation. To address the challenge of scene perception, our perception module integrates advanced tools such as Recognize-Anything [6], Grounding DINO [7], Segment Anything-2 [8], and CLIP [9] to detect objects or components and extract their semantic features. In the memory module, each object within the scene is represented as a node characterized by geometric and semantic features, with the relationships between objects encoded in the graph’s edges. The module continuously updates object features and scene graphs by locally refining the areas where the robot interacts, preserving knowledge to support ongoing exploration and exploitation, while avoiding the need to reconstruct the entire scene and enabling more efficient adaptation to dynamic environments.
Our task planning module uses the advanced large language models to decompose tasks into manageable subtasks. Then, the navigation and manipulation module are activated to execute the planned actions. In experiments, we deployed a mobile robot platform in multiple real-world indoor environments, conducting long-term task trials under varying degrees of manual environmental modifications. We demonstrated that DovSG can accurately update the 3D scene graph and perform excellently in long-term tasks and subtasks such as pick-up, place, and navigation. Our contributions are as follows:
@@ II Related Works
3D scene representation in robotics often follows two main approaches: using foundation models to create 3D structures [10, 11, 12], or combining 2D image and vision-language models [7, 9, 8, 6] to connect with the 3D world, showing impressive results in open-vocabulary tasks [13, 14], enabling language-guided object grounding [12, 3, 13] and 3D reasoning [15, 16, 17]. However, dense semantic features for each point are redundant, memory-intensive, and not easily decomposable, limiting their use in dynamic robotics applications.
@@
3D scene graphs use a hierarchical graph structure to represent objects as nodes and their relationships as edges within a scene [18]. This approach efficiently stores semantic information in robot memory, even in large and dynamic environments. Recent methods like ConceptGraphs [13] and HOV-SG [19] create more compact and efficient scene graphs by merging features of the same object across multiple views, thereby reducing memory redundancy. However, these methods often assume static environments and overlook scene updates during interactive robot manipulations. RoboEXP [14] addresses this by encoding both spatial relationships and logical associations that reflect the effects of robot actions, enabling the discovery of objects through interaction and supporting dynamic scene updates. Building on this, we have applied these concepts to mobile robots, allowing them to rapidly update memory and adjust the 3D scene graph in real time during dynamic interactions with environments.
@@ IV-2 Environment and Task Setups
To verify our method’s ability to enable robots to perform long-term tasks in dynamic environments, we designed an experiment in 4 real-world rooms. The experiment simulated dynamic environments through human interactions, which modified object positions, added new objects, or revealed hidden ones. Each experiment began with a 3D scene graph and involved two consecutive tasks. Before the first task, we manually adjusted objects related to the second task by changing their positions, adding new objects, or revealing hidden ones. These modifications were designed to be detectable during the execution of the first task. After completing the first task, instead of being repositioned to an initial start point, the robot continued to the next task from its current state. This setup simulated real-world scenarios with continuous environmental changes caused by human interference. To evaluate the robot’s ability to detect and adapt to environmental changes, we categorized the modifications to the objects in the second task into three levels (see Fig. 4):
@@ IV-5 Results
TABLE III: Effieciency comparison between Ok-Robot and DovSG
@@
V Conclusion
@@
Acknowledgement
This project funded by the grant from the National Natural Science Foundation of China (No. 52205244) and supported by the Ministry of Industry and Information Technology (MIIT) Key Laboratory of Intelligent Manufacturing for High-end Aerospace Products, and the Beijing Key Laboratory of Digital Design and Manufacturing.

### v5

@@ I Introduction
Mobile manipulation in real-world environments is increasingly expected to perform complex, long-term tasks that require adaptability and resilience to changes. These dynamic environments are characterized by frequent alterations caused by human activities, robot interactions, or the inherent variability of the surroundings. Traditional robotic systems often fall short in such settings because they typically rely on the assumption of a static or minimally changing environment [1, 2, 3, 4, 5]. This limitation restricts their applicability in real-world scenarios where adaptability is crucial.
In this work, we enhance robotic capabilities by introducing a novel and practical robotic framework, the DovSG system. This framework comprises five key modules: perception, memory, task planning, navigation, and manipulation, as illustrated in Fig. Dynamic Open-Vocabulary 3D Scene Graphs for Long-term Language-Guided Mobile Manipulation. To address the challenge of scene perception, our perception module integrates advanced tools such as Recognize-Anything [6], Grounding DINO [7], Segment Anything-2 [8], and CLIP [9] to detect objects or components and extract their semantic features. In the memory module, each object within the scene is represented as a node characterized by geometric and semantic features, with the relationships between objects encoded in the graph’s edges. The module continuously updates object features and scene graphs by locally refining the areas where the robot interacts, preserving knowledge to support ongoing exploration and exploitation, while avoiding the need to reconstruct the entire scene and enabling more efficient adaptation to dynamic environments. Our task planning module uses the advanced large language models to decompose tasks into manageable subtasks. Then, the navigation and manipulation module are activated to execute the planned actions. In experiments, we deployed a mobile robot platform in multiple real-world indoor environments, conducting long-term task trials under varying degrees of manual environmental modifications. We demonstrated that DovSG can accurately update the 3D scene graph and perform excellently in long-term tasks and subtasks such as pick-up, place, and navigation. Our contributions are as follows:
@@ II Related Works
3D scene representation in robotics often follows two main approaches: using foundation models to create 3D structures [10, 11, 12], or combining 2D image and vision-language models [7, 9, 8, 6, 13] to connect with the 3D world, showing impressive results in open-vocabulary tasks [14, 15], enabling language-guided object grounding [12, 3, 16, 14] and 3D reasoning [17, 18, 19]. However, dense semantic features for each point are redundant, memory-intensive, and not easily decomposable, limiting their use in dynamic robotics applications.
@@
3D scene graphs use a hierarchical graph structure to represent objects as nodes and their relationships as edges within a scene [20]. This approach efficiently stores semantic information in robot memory, even in large and dynamic environments. Recent methods like ConceptGraphs [14] and HOV-SG [21] create more compact and efficient scene graphs by merging features of the same object across multiple views, thereby reducing memory redundancy. However, these methods often assume static environments and overlook scene updates during interactive robot manipulations. RoboEXP [15] addresses this by encoding both spatial relationships and logical associations that reflect the effects of robot actions, enabling the discovery of objects through interaction and supporting dynamic scene updates. Building on this, we have applied these concepts to mobile robots, allowing them to rapidly update memory and adjust the 3D scene graph in real-time during dynamic interactions with environments.
@@ IV-2 Environment and Task Setups
To verify our method’s ability to enable robots to perform long-term tasks in dynamic environments, we designed an experiment in 4 real-world rooms. The experiment simulated dynamic environments through human interactions, which modified object positions, added new objects, or revealed hidden ones. Each experiment began with a 3D scene graph and involved two consecutive tasks. Before the first task, we manually adjusted objects related to the second task by changing their positions, adding new objects, or revealing hidden ones. These modifications were designed to be detectable during the execution of the first task. After completing the first task, instead of being repositioned to an initial start point, the robot continued to the next task from its current state. This setup simulated real-world scenarios with continuous environmental changes caused by human interference.
To evaluate the robot’s ability to detect and adapt to environmental changes, we categorized the modifications to the objects in the second task into three levels (see Fig. 4):
@@ IV-5 Results
TABLE III: Effieciency comparison.
@@
V Limitation and Future Work
Limitation: The performance of DovSG’s visual relocalization depends on the presence of distinctive visual cues (e.g., textures, edges, objects) for key point matching. In environments with sparse, minimal, or repetitive features, the system may struggle to extract enough distinctive points, impacting localization accuracy.
Future work: Exploring real-time precise localization through multi-sensor fusion, efficient memory update mechanisms, and developing lightweight methods for scene representation is crucial. Additionally, enabling efficient collaboration between mobile robots and manipulators will be key to making mobile manipulation more practical in real-world scenarios.
VI Conclusion